OpenStudy (anonymous):

will fan and medal! find the function rule. x -2 -1 0 1 2 y 94 -1 -6 -11

OpenStudy (michele_laino):

we are searching for a function like this: \[y = ax + b\] where \(a,b\) are the coefficient to be established, and \(x,y\) are the unknowns

OpenStudy (anonymous):

okay

OpenStudy (michele_laino):

let's consider the first pair, namely \(x=-2,y=9\) substituting those values I get: \(-2a+b=9\) am I right?

OpenStudy (anonymous):

I think so

OpenStudy (michele_laino):

please try to substitute \(x=-1,y=4\) into my equation \(y=ax+b\), what do you get?

OpenStudy (michele_laino):



OpenStudy (anonymous):

I don't know

OpenStudy (michele_laino):

we get this equation: please look at my hint: \(4=-a+b\)

OpenStudy (anonymous):

will b = -6?

OpenStudy (michele_laino):

I got \(b=-1\)

OpenStudy (michele_laino):

we have to solve this algebraic system: \[\left\{ \begin{gathered} - a + b = 4 \hfill \\ - 2a + b = 9 \hfill \\ \end{gathered} \right.\]

OpenStudy (anonymous):

I don't get this

OpenStudy (michele_laino):

when I replaced x with -2 and y with 9 I got this equation: \[ - 2a + b = 9\] when I replaced x with -1 and y with 4 I got this equation: \[ - a + b = 4\] so, in order to compute the values of \(a,b\) we have to solve this system: \[\left\{ \begin{gathered} - a + b = 4 \hfill \\ - 2a + b = 9 \hfill \\ \end{gathered} \right.\]

OpenStudy (anonymous):

I know its against the rules but this is a timed test can I please get a direct answer

OpenStudy (michele_laino):

no, I'm sorry, I can not give the direct answer, since it is against the Code of Conduct

OpenStudy (anonymous):

ugh

OpenStudy (anonymous):

poop

OpenStudy (michele_laino):

please solve the subsequent equation, with respect to \(b\): \[ - a + b = 4\] what do you get?

OpenStudy (anonymous):

I don't know but I swear if I can just get a direct answer I will fan and medal you, you will get the reward you deserve

OpenStudy (anonymous):

What is "x -2 -1 0 1 2 y 94 -1 -6 -11" even supposed to be?